OpenStudy (anonymous):

A firecracker shoots up from a hill 165 feet high with an initial speed of 110 feet per second. Using the formula H(t) = −16t^2 + vt + s, determine how long it will take the firecracker to hit the ground.

OpenStudy (nathan917):

We have v = 105 ft/s and s = 160 ft. Sub those in: H(t) = -16t^2 + 105t + 160 We want to know when this hits the ground, so set H(t) equal to zero: 0 = -16t^2 + 105t + 160 Use the quadratic formula: t = [1 / (2 * -16)][-105 ± √((105)^2 - 4(-16)(160))] t = (105 / 32) ± (1 / 32)√(21265) So we have two roots, but we obviously want the positive one, because negative time makes no sense. Thus: t = (105 / 32) + (1 / 32)√(21265) s t ≈ 7.84 s The answer is B.

OpenStudy (anonymous):

howd u know it was b? lol u taking conections?

OpenStudy (nathan917):

Yes. but I'm Not In Your grade :P 7.5 seconds 7.8 seconds 8.1 seconds 8.3 seconds These you options? :P

OpenStudy (anonymous):

it wasnt in that oreder but 7.8 was there so i got it it was c. on mine lol

OpenStudy (nathan917):

:P. They change it every year or so.

OpenStudy (anonymous):

oh lol ... well thnx :)